Identify the zeros of f(x) = (x - 7)(x + 4)(3x - 2). 4, ,7 0 -4, 2 3 7 O-4, WIN O-4, -7
Answer #1

The given fonction fcn) = (x-7)(x+4) (3x-2) For the zeros of fead, fla=0 (x-7)(x+4) (3x-2) = 0 7-7= 0 Now or at 4-0 or 32e-2=